# A Closed-Form Solution for Kernel Adaptive Filtering

### v2

@@ 4.1 Demonstration of the Difference Equation Interpretation
We will now use the interpretation given in section 3.5 to discover the underlying functions that generate a time series. First, we will discuss the system used in this experiment. The input r.p., $X$, is white Gaussian noise drawn from the distribution $\mathcal{N}(0,\pi)$. The model output, $Z$, is created by applying the nonlinear mapping given in (40) to $X$.
@@
$\begin{split}Z_{t}=&0.5\tanh(X_{t})^{2}+\sin(X_{t-1})^{3}+0.5\tanh(X_{t-2})^{3}\\ &+0.2\sin(X_{t-3})^{2}+0.75\tanh(X_{t-4})^{2}\end{split}$ (40)
The FWF is then tasked with finding a mapping function from $X$ to $Z$ i.e., a conventional system identification setting. The bottom half of figure 3 shows the $L$ functions (one for each value of $\tau$) found by the FWF. The green histograms show the p.d.f of the realizations used in the training set, scaled for visualization. We observe that in regions covered in the training set, the functions learned by the FWF are vertically shifted (biased) versions of the true functions given in equation (40). However, when these biased versions are summed together, they converge to the true difference equation given in (40). Centering the covariance in RKHS will compensate for the bias. It’s also interesting that the FWF outputs are constant for values of $\tau$ which are larger than the true memory depth of the system.
The top right plot in Figure 3 compares the performance of the FWF with the other nonlinear filtering methods mentioned above. Each method was tested with 5 different training and testing windows at each value of $N$. The average test set MSE is shown with error bars, giving the variance of the test set MSE across the different training and testing windows. For each method, a search across a range of kernel sizes (0.7520.75-2 at increments of $0.1$) was performed, and the best results for each method are shown. The FWF provides a clear performance boost over the other methods on this task. This suggests that if the model assumptions hold, then its performance is better than the other nonlinear filtering methods. Moreover, we can plot the underlying functions learned by the FWF at each lag and can interpret the results as a difference equation.
@@ 4.2 Mackey-Glass Prediction
The Mackey-Glass is a chaotic nonlinear time series with equation
$\frac{dx_{t}}{dt}=\frac{a\theta^{n}x_{t-\tau}}{\theta^{n}+x_{t-\tau}^{n}}-bx_{t}$ (41)
This time series was introduced in [21] as a model capable of producing nonlinear chaotic behavior similar to respiratory and hematopoietic diseases. It is commonly used to test time series prediction methods (see [7, 19, 17]). In this experiment we use the values $a=0.2$, $b=0.1$, $n=10$, $\theta=1$ and $\tau=30$. The time series is then discretized with a sampling period of 6 seconds.
Figure 4: Comparison of Test MSE on Mackey-Glass task window sizes of L=20L=20(left) and L=15L=15(right).
Figure 5: Comparison of Test MSE on noisy Mackey-Glass with window sizes of L=20L=20(left) and L=15L=15(right), $N=2000$ for all noise levels.
A Gaussian kernel was used for all kernel methods. Figure 4 shows the test set MSE as a function of the number of training samples ($N$) for window sizes ($L$) of 15 and 20. For each value of $N$ we use 5 different training windows and take the average test MSE across these trials. We perform a hyperparameter search for each method and show the best MSE for each method in Figure 4. We tested the FWF with sample embedding sizes($D$) of 1,2, and 3. For FWF with sample embeddings of 2 and 3, we use a truncation point ($K$) of 10. For the FWF with $D=1$, we use a truncation point of 25. Larger values of $K$ were tested, but did not improve performance. For this task, increasing $D$ improves the performance of the FWF markedly. When compared to the other kernel methods, the FWF with $D=2,3$ is on par with KRLS and GPR for $L=15$, and outperforms KRLS and GPR for $L=20$. For the FWF with $D=1$, the MSE flattens at 500 training samples. For the FWF with $D=2,3$ more samples are needed for the MSE curve to flatten, which supports our claim of the trade-off between more general spaces and the amount of data necessary to estimate the MMSE solution.
We also tested robustness to input noise by adding white Gaussian noise at increasing levels to the input of the model. Figure 5 shows the performance for each method for a given signal-to-noise ratio. The same hyperparameter search done for the noiseless prediction case was used. We see that the FWF exhibits similar sensitivity to noise as the other methods.
@@ 4.3 Lorenz Attractor
$\frac{dx}{dt}=\sigma(y-x),\quad\frac{dy}{dt}=x(\rho-z)-y,\quad\frac{dz}{dt}=xy-\beta z$ (42)
@@
4.4 SunSpot
In this experiment, we test the methods on a real-world dataset (measured rather than simulated), the sunspot data [8]. This time series contains monthly averages of the daily sunspot numbers reported from the WDC-SILSO, Royal Observatory of Belgium. The time series is standardized to have zero mean and a standard deviation of 1. The task for the models is to forecast the number of sunspots 10 samples in the future. A grid search (0.7530.75-3 in increments of $0.25$) for the best kernel size for each kernel method was performed. The number of training samples is 2000 and the test set is 300 samples. The average MSE and standard deviation for the best set of hyperparameters across 5 different training windows are given in 3. $L=10$ for all methods. On this task, the FWF with $D=2$ performs best in the test set.
@@ 5 Conclusions
In summary, we have successfully extended Parzen’s MMSE in $\mathcal{H}_{R}$, to a nonlinear data-dependent RKHS, $\mathcal{H}_{U}$. In $\mathcal{H}_{U}$, the MMSE solution can be expressed as the cross-covariance function. Essentially, when the statistics of the input r.p. are embedded into the inner product of a space, the orthogonal projection of any desired r.v. (i.e. the MMSE solution) is immediately given by the cross-covariance function.
The calculation of the nonlinear MMSE solution is simplified with the use of an explicit finite-dimensional RKHS that approximates the universal Gaussian kernel. In this finite-dimensional space, the $U$ operator takes the form of a finite-dimensional matrix, which is used to define the inner product in $\mathcal{H}_{\bm{U}}$. This is a strength of using a finite-dimensional RKHS, as it greatly simplifies the calculation of inner products in $\mathcal{H}_{U}$.
Although the FWF belongs to the KAF family, its principles are different and represent a fresh approach to exploit the linear structure of the Hilbert space. Rather than implementing search techniques in a data-independent space, we focus on first building the RKHS in which we define our solution. Then show that search is not necessary in this space. Another distinction between FWF and other KAF is that the FWF solution defines functionals over time and space (i.e. the amplitude values of the signal).
In future work, we should address some practical limitations of the method. One of these limitations is that the dimensionality of $\mathcal{H}_{U}$ grows exponentially with sample embedding size. Other methods for explicit space approximations, which are more efficient in terms of dimensionality, are needed. We should also further investigate how sensitive the solution is to data non-stationarity and seek methods for lifting this assumption. We should also expand on the interpretation of the FWF as a nonlinear difference equation and see what sorts of insights can be gained from this interpretation. Finally, one deeper question stemming from this work is: Can we define a practical technique for taking inner products in $\mathcal{H}_{U}$ without first referencing an extrinsic coordinate system? Answering this question would yield a completely data-determined Hilbert space which is truly the “natural” setting for taking conditional expectations with respect to $X$.
@@ Appendix A The $\mathbf{U}$ matrix
In section 3.2 an abbreviated construction of the $\mathbf{U}$ matrix is given. Here we show a more explicit construction of this matrix as a block matrix. Due to strict stationarity, we can remove the dependence on $s,t$ and let $s=t-\tau$. Then
$\bm{U}(\tau,n,m)=\mathbb{E}[\phi_{n}(\bm{X}_{t}^{D})\phi_{m}(\bm{X}_{t-\tau}^{D})]$ (43)
Intuitively, $\bm{U}(\tau,n,m)$ measures the correlation between the projection of $\bm{X}_{t}^{D}$ and $\bm{X}_{t-\tau}^{D}$ across all dimensions of $\mathcal{H}_{S}$. For each value of $\tau=0,1,\dots,L-1$, we will get a $M\times M$ dimensional matrix,
$\mathbf{U}_{\tau}=\begin{bmatrix}\bm{U}(\tau,1,1)&...&...&\bm{U}(\tau,1,M)\\ \vdots&\ddots&...&\vdots\\ \bm{U}(\tau,M,1)&...&...&\bm{U}(\tau,M,M)\end{bmatrix}=\mathbb{E}[\phi(\bm{X}_{t}^{D})\phi(\bm{X}^{D}_{t-\tau})^{\top}]$ (44)
Under the assumption of finite model order, the matrix $\mathbf{U}$ is a $((M\cdot L)\times(M\cdot L))$ dimensional positive semi-definite matrix,
@@
Since $\mathbf{U}$ is the covariance matrix of $\phi(\bm{X}_{t}^{DL})$ (the same for all $t$ because of stationarity), it is a positive semi-definite matrix [24]. The $\mathbf{U}$ matrix describes the correlations of the projection of $\bm{X}^{DL}$ across both space (the features of $\mathcal{H}_{S}$) and time.
Now we show exactly how $\mathbf{U}$ is related to $U(s,t,x,y)$. Assuming strict stationarity, $U(s,t,x,y)$ can be written as a function of $\tau=|s-t|$. The following relationship between the submatrices of $\mathbf{U}$ and $U(\tau,x,y)$ is given as,
$U(\tau,x,y)\approx\sum_{n=1}^{M}\sum_{m=1}^{M}\bm{U}(\tau,n,m)\phi_{n}(x)\phi_{m}(y)$ (46)
where $\bm{U}(\tau,n,m)$ is defined in equation (43). A simulation that empirically validates 46 is given in the supplementary materials.
Funding
This material is based upon work supported by the Office of the Under Secretary of Defense for Research and Engineering under award number FA9550-21-1-0227, and partially supported by ONR grants N00014-21-1-2295 and N00014-21-1-2345. This research was supported by the Science, Mathematics and Research for Transformation (SMART) Scholarship Program.

### v3

@@ 4.1 Demonstration of the Difference Equation Interpretation
The interpretation given in section 3.4 is now applied to identify the underlying functions that generate a time series. The simulated data is generated as follows. The input to the system ($x$) is white Gaussian noise with i.i.d. samples drawn from the distribution $\mathcal{N}(0,\pi)$. The system output, $z$, is obtained via the nonlinear mapping
$\small\begin{split}z_{t}=&0.5\tanh(x_{t})^{2}+\sin(x_{t-1})^{3}+0.5\tanh(x_{t-2})^{3}\\ &+0.2\sin(x_{t-3})^{2}+0.75\tanh(x_{t-4})^{2}.\end{split}$ (19)
@@
The task is to estimate the mapping from $x$ to $z$. Note this is a standard setup for system identification.
The bottom half of Figure 3 shows the set of functions $\{f^{*}_{\tau}(\cdot)\}^{L-1}_{\tau=0}$ found by the FWF. The green histograms show the p.d.f of the input signal used for training, scaled for visualization. We observe that in regions covered in the training set, the functions learned by the FWF are biased versions of the true functions given in equation (19). However, when these biased versions are summed together, they converge to the true difference equation given in (19). Centering the covariance in RKHS will compensate for the bias. It is noteworthy that the FWF outputs remain constant for values of $\tau$ exceeding the true memory depth of the system.
The top right plot in Figure 3 compares the performance of the FWF with the other methods mentioned above. Each method was tested with five independent training and testing windows at each value of $N$. The average test set MSE is shown. The FWF provides a clear performance boost over the other methods on this task. This suggests that if the model assumptions hold, then the FWF’s performance is better than the other nonlinear filtering methods. Moreover, we can plot the underlying functions learned by the FWF at each lag and can interpret the results as a difference equation.
@@ 4.2 Mackey-Glass Prediction
The Mackey-Glass time series is a chaotic nonlinear time series governed by the equation,
$\frac{dx_{t}}{dt}=\frac{a\theta^{n}x_{t-\tau}}{\theta^{n}+x_{t-\tau}^{n}}-bx_{t}.$ (20)
This time series was introduced in [38] as a model capable of producing nonlinear chaotic behavior similar to respiratory and hematopoietic diseases, and is commonly used to test time series prediction methods (see [6, 2]). We use the values $a=0.2$, $b=0.1$, $n=10$, $\theta=1$ and $\tau=30$. The time series is then discretized with a sampling period of 6 seconds using the fourth-order Runge-Kutta method.
Figure 4: (top)Comparison of Test MSE vs number of training samples(N) Mackey-Glass with window sizes of L=20L=20(left) and L=15L=15(right). (bottom)Comparison of Test MSE on noisy Mackey-Glass with window sizes of L=20L=20(left) and L=15L=15(right), $N=2000$ for all SNR levels. Error bars indicate best and worst case performance.
The top two plots in Figure 4 show the test set MSE as a function of the number of training samples ($N$) for window sizes $L=15,20$. For each value of $N$ we give the average, best case, and worst case test MSE across five independent training and testing windows. We tested the FWF with sample embedding sizes of $D=1,2,3$. Increasing $D$ consistently improves the performance of the FWF. When compared to the other kernel methods, the FWF with $D=2,3$ is on par with KRLS and GPR for $L=15$, and outperforms KRLS and GPR for $L=20$. The FWF with $D=1$ plateaued after 500 samples, whereas larger sample embedding sizes required more data to converge. This supports our assertion of the trade-off between model capacity and data requirements for the estimation of the FWF solution.
To assess robustness, we compare test MSE when additive white Gaussian noise at varying signal-to-noise ratios (SNR) corrupts the input to the model. The bottom two plots in 4 show the performance for each method at each SNR.
@@ 4.3 Lorenz Attractor
$\frac{dx}{dt}=\sigma(y-x),\qquad\frac{dy}{dt}=x(\rho-z)-y,\qquad\frac{dz}{dt}=xy-\beta z$ (21)
@@
4.4 Sunspot Forecasting
In this experiment, we test the methods on a real-world dataset (measured rather than simulated). The sunspot data [42] contains monthly averages of the daily sunspot numbers reported from the WDC-SILSO, Royal Observatory of Belgium. The time series is standardized to have a mean of zero and a standard deviation of 1. The task for the models is to forecast the number of sunspots 10 samples in the future. The number of training samples is 2000, the number of test set samples is 300, and $L=10$ for all methods. The average MSE and standard deviation across 5 independent training and testing windows are given in Table 3. While the FWF with $D=3$ outperformed the other instantiations of the FWF in the previous experiments, on this task, it exhibited overfitting and large variance in test set performance, likely due to the increase in noise in this dataset.
@@ 5 Conclusions
In summary, we have successfully extended Parzen’s MMSE in $\mathcal{H}_{R}$, to a nonlinear data-dependent RKHS, $\mathcal{H}_{U}$. By embedding the input random process statistics into the inner product of the space, the orthogonal projection of any desired r.v. (the MMSE solution) is immediately given by the cross-covariance function. Rather than implementing search techniques in a data-independent space, we build the RKHS in which our solution is defined, rendering explicit search unnecessary. Calculation of the FWF is simplified using an explicit finite-dimensional RKHS approximating the universal Gaussian kernel, where the $U$ operator becomes a finite-dimensional matrix defining the inner product in $\mathcal{H}_{\bm{U}}$. Experimentally, we demonstrate that when the assumptions made by the FWF are met, we outperform other kernel-based nonlinear adaptive filtering methods. Moreover, we can interpret the FWF solution as a nonlinear difference equation and extract the underlying functions learned by the FWF.
Future work should address some practical limitations of the method. One of these limitations is that the dimensionality of $\mathcal{H}_{\bm{U}}$ grows exponentially with sample embedding size and truncation point. Other methods for explicit space approximations, perhaps ones that are more efficient in terms of dimensionality, should be explored. We should also expand on the interpretation of the FWF as a nonlinear difference equation, as this is unique to the FWF. Finally, one deeper question stemming from this work is: Can we define a practical technique for taking inner products in $\mathcal{H}_{U}$ without first referencing an extrinsic coordinate system? Achieving this would yield a fully data-determined Hilbert space which is truly the “natural” setting for conditional expectations with respect to $X$.
Funding and AI Use
This material is based upon work supported by the Office of the Under Secretary of Defense for Research and Engineering under award number FA9550-21-1-0227, and partially supported by ONR grants N00014-21-1-2295 and N00014-21-1-2345. This research was supported by the SMART Scholarship Program. The authors used ChatGPT to identify spelling, grammar, and tense consistency issues. Each of these suggestions was reviewed by the authors, and the final edit was made.
@@ Appendix A The $\mathbf{U}$ matrix
In section 3.2 an abbreviated construction of the $\mathbf{U}$ matrix is given. Here we show a more explicit construction of this matrix as a block matrix. Due to strict stationarity, we can remove the dependence on $s,t$ and let $s=t-\tau$. Then $\bm{U}(\tau,n,m)=\mathbb{E}[\phi_{n}(\bm{X}_{t}^{D})\phi_{m}(\bm{X}_{t-\tau}^{D})]$. Intuitively, $\bm{U}(\tau,n,m)$ measures the correlation between the projection of $\bm{X}_{t}^{D}$ and $\bm{X}_{t-\tau}^{D}$ across all dimensions of $\mathcal{H}_{S}$. For each value of $\tau=0,1,\dots,L-1$, we can store $\bm{U}(\tau,n,m)$ in a $M\times M$ dimensional matrix, $\bm{U}_{\tau}=\mathbb{E}[\phi(\bm{X}_{t}^{D})\phi(\bm{X}^{D}_{t-\tau})^{\top}]$. Finally, the matrix $\mathbf{U}$ is a $((M\cdot L)\times(M\cdot L))$ dimensional positive semi-definite matrix,
@@
Since $\mathbf{U}$ is the covariance matrix of $\phi(\bm{X}_{t}^{DL})$ (the same for all $t$ because of stationarity), it is a positive semi-definite matrix [43]. The $\mathbf{U}$ matrix describes the correlations of the projection of $\bm{X}^{DL}$ across both space (the features of $\mathcal{H}_{S}$) and time. Assuming strict stationarity, the following relationship between the elements of the submatrices of $\mathbf{U}$ and $U(\tau,x,y)$ is given as,
$U(\tau,x,y)\approx\sum_{n=1}^{M}\sum_{m=1}^{M}\bm{U}(\tau,n,m)\phi_{n}(x)\phi_{m}(y).$ (23)